# Inferring Capabilities from Task Performance with Bayesian Triangulation

### v1

@@ 5.1 The Testbed and Synthetic Agents
Team A generated 30 synthetic agents of four different types to ensure a diversity of behaviour. Reference Agents act as reference points, e.g., a perfect agent, an agent with $x$success, etc.
@@
Fraudster Agents behave in a sophisticated manner that might look like object permanence, but actually is not (Uzgiris and Hunt 1975). Tables  6 and  7 in appendix D.4 show the specifics of each agent. This information was not disclosed to Team B.
5.2 Measurement Layout
Object permanence is a difficult capability to evaluate—combined with the need to visually identify goals, navigate to the goal, avoid hazards, and to remember the location of the object; there are many possible factors influencing overall performance that need to be disambiguated.
We present the measurement layout for the OP tasks in Fig. 5. The same broad structure as for the AAIO measurement layout is expanded upon. In the OP task-suite, there are additional challenges to navigation in the form of ramps, platforms, and lava. Subsequently, nodes for the agent’s capability (and instance-level performance) at handling each of these must be added to the layout. Meta-features for the presence of each of these challenges are needed. To represent object permanence specifically, we add extra meta-features: Occluder Presence, Time Under Occlusion, and Number of Positions (denoting the number of search-location choices available to the agent in forced choice instances). We further add capability nodes for Object Permanence and Memory Ability (each with an associated Performance node). Memory Performance, Occluder Presence and Number of Positions all feed into object permanence performance. The linking function for OP performance is the non-complementary product of Memory Performance and the logistic of the OP margin. The OP margin is the difference between Object Permanence and the product of Number of Positions and Time Under Occlusion. If no occluder was present the linking function for Object Permanence Performance returns $1$.
5.3 Inference and Analysis
Team B took the synthetic agent results and inferred the cognitive profiles using the PyMC implementation of the measurement layout. Overall, our model was successful at inferring system capabilities from patterns of performance. We evaluated the model in three ways. The first was a qualitative comparison of the cognitive profile descriptions generated by Team B with the criteria used to generate the agents. The model facilitated a correct description of the agent in around 75% of cases. For the “Achilles Heel” agents in particular, which had object permanence abilities but a specific weakness, the model was highly successful at identifying the cause of failure. It was, however unable to spot “fraudsters” making use of strategies (such as “go to the last seen location of the reward”) to mimic OP capability. Appendix D.4 contains a detailed breakdown of inferred parameters of the cognitive profiles of each agent, more information on the experimental set up, and the full qualitative analysis comparing inferred capabilities to the ground truth capabilities of the agents designed by Team A.
Capability OP FlatNav Visual Lava Platform Ramp Memory
Overall RMSE 0.28 0.18 0.16 0.2 0.28 0.38 0.15
Included RMSE 0.13 0.11 0.24 0.17 0.134 0.27 0.16
(a) AAIO Tasks
(b) O-PIAAGETS tasks.
Figure 6: Brier score of model predictions against agent success rates.
The second validation involved calculating the RMSE of the model for each ability (normalised to be in $[0,1]$), relative to the ground truth generated by Team A. Given in Table 1, model errors were modest across most abilities. Understandably, the model struggled most with agents that were created using features not captured by the model. For example, agent 23 was generated to have robust object permanence only in cup tasks (passing only 10% of other types). As our model did not include task type as a meta-feature, it erroneously took the non-cup task failures as evidence the agent lacks OP as well as struggling with platforms. However, given that the cup tasks are the only OP tasks without platforms, this is not an unreasonable assessment. By contrast, as the bottom row of the table shows, the model was accurate for agents generated using only features included in the model. This highlights the importance of having a measurement layout capturing the important meta-features that most affect performance, and also of designing test-batteries that allow confounding factors to be disentangled.
For our third validation, we show that our model can be used for accurate prediction of agent success on unseen task instances. Taking instance meta-features and applying forward inference with our model yields a probability of success. We compare this to the probability predicted by aggregate success using a Brier score. The full results are given in Appendix E, but a summary is given in Figure 5(b). The model is consistently a better predictor of success than relying on the aggregate measure. We see better prediction performance for agents with less certain success rates. These are the agents for which we are more interested in improving prediction as there is more capacity to reduce uncertainty.
6 Discussion
Increasingly more machine learning systems are (pre-)trained once and deployed for a range of tasks, Task-oriented evaluation based on aggregated performance cannot robustly evaluate these systems across the many situations in which they might be deployed, nor predict behaviour for new tasks and changing distributions. However, we can often identify abstract capabilities that are not directly measurable, such as ‘object permanence’ in agents or ‘handling negation’ in language models. In the behavioural sciences this is a common approach, but many techniques are populational. Triangulation for single individuals by contrasting against the task demands is common in animal cognition research, but it is unusual to see complex inferences from thousands of items, as is common in machine learning. In this paper we have integrated several elements to make this inferential exercise possible and shown it can identify the capabilities and biases of each model—its cognitive profile—independently.
This unleashes a new way of evaluating general-purpose systems, by identifying areas of weakness or reasons for task failure. This is especially useful for hierarchical capabilities with multiple levels of dependency. We can speed up the debugging process by isolating skill-gaps or biases responsible for failure, and improve confidence in predictions of performance on new tasks. Instead of hoping that the average performance will hold on new tasks, our approach enables the use of task demands and the inferred cognitive profile to determine whether system deployment is safe and worthwhile.
There are several limitations of this work. If many parameters must be inferred simultaneously then many evaluation instances are needed. We advocate incremental approaches, where each capability is inferred with subsets of tasks before moving to tasks with more complex dependencies. Nonetheless, we have shown that even a posteriori, as in the AAIO, inferences can be achieved simultaneously. Here we covered only the domain of agentic systems, but the same ideas can be extrapolated to language, vision and other domains. Ultimately, the biggest challenge for the construction of measurement layouts is the need for detailed domain knowledge and robust evaluation batteries. Both can be costly, but are urgently needed to improve AI evaluation. Currently, evaluation is focused on advancing SOTA performance on an increasingly small selection of benchmarks that often lack construct validity and are limited in scope (Koch et al. 2021; Raji et al. 2021).
Our approach also highlights the need for a clear distinction between training datasets, which can be massive, unannotated, and built in an adversarial way (Kiela et al. 2021) (Jia and Liang 2017; Zellers et al. 2019), and evaluation batteries, which should be devised carefully with systematic variation across the elements that might affect performance. Our vision is that, from well-constructed batteries such as O-PIAAGETS, building measurement layouts and triangulating with them should become routine.
Appendices
Here we include some supplementary material to the paper. First we include more details about PyMC, the Bayesian inference engine we are using and the decisions made for the estimation algorithm, and then we give more details and results for the two scenarios; first, the Animal AI Olympics and second, O-PIAAGETS.
Appendix A PyMC
PyMC (Salvatier, Wiecki, and Fonnesbeck 2016) is a probabilistic programming language for building Bayesian models and fitting them with Markov Chain Monte Carlo (MCMC) methods (Andrieu et al. 2003). Complex Bayesian models often require solving intractable integrals for inference such as the normalisation constant in Bayes’ theorem,
[…]
among others. Monte Carlo (MC) Integration provides a method to approximate these intractable integrals using random sampling.
In order to evaluate
[…]
we can rewrite it as an expectation
[…]
for any density function $p$ such that $p(x)>0$.
MC Integration takes a random sample $X=\{x_{i}\}_{i=1}^{N}\in\Omega$ with distribution $p(x)$ and uses the Law of Large Numbers to approximate the integral as:
$I\approx\frac{1}{N}\sum_{i=1}^{N}\frac{f(x_{i})}{p(x_{i})}$
MC Integration is reliant on being able to sample efficiently from $\Omega$ according to $p(x)$. MCMC algorithms aim to do this, scaling to complex, hierarchical models by building a Markov Chain with equilibrium distributions approximating the desired probability density function $p$. Sampling from the Markov Chain can then be used as a stand-in for sampling $p(x)$. The prototypical example of an MCMC algorithm is the Metropolis Hastings algorithm (Hastings 1970). Hamiltonian Monte Carlo methods (Betancourt and Girolami 2013) improve on typical MCMC algorithms by replacing the random-walk element exploring the Markov Chain state-space with Hamiltonian Mechanics.
PyMC requires all expressions to be differentiable, which means that functions such as maximum and minimum are implemented as generalised means, and logical operators are implemented with products and sums.
A.1 Measurement Layouts in PyMC
In our work, we utilise the No U-Turn Sampler (Hoffman and Gelman 2014) built into PyMC, an extension to Hamiltonian MC methods which eliminates the need to set particular hyper-parameters.
With the NUTS sampler we use 2 sampling chains, tuned for $1000$ samples that are discarded in order to optimise the step-size within NUTS, and then draw $1000$ samples to estimate the posterior.
We define the model once for each scenario and then we estimate the parameters of the cognitive profile and some other INON elements for each system being evaluated independently.
As mentioned in the introduction, data and code will become available upon publication.
Appendix B Animal AI Environment and Competition
Animal-AI is an environment for training and testing AI systems. It is built in the Unity engine (Juliani et al. 2018). Animal-AI provides a first-person perspective in a 3D environment. The environment can contain a number of pre-specified objects, representing rewards, obstacles, tools, or dangers. Tasks can be defined using a domain-specific language to procedurally generate a large number of tasks instances. The agent’s interaction with the world is limited to moving around in the 3D space, though it can exploit physics interactions to exert more influence over the environment. Fig. 7 shows two screenshots from separate task instances.
In the Animal-AI Olympics (Crosby et al. 2020), an example arena containing the possible objects was provided to contestants. They were instructed that their agents would be evaluated on a variety of problems constructed from the objects in the example arena. The contestants then needed to build an agent—using any method of their choosing—that they believed would succeed at a wide variety of tasks. The tasks that ultimately made up the evaluation suite were based on adaptations of tasks from comparative and developmental literature. More information can be found at (http://animalai.org/AAI/).
For our work with the AAIO data-set we utilised a subset of tasks instances that were solvable with simple goal-directed behaviour. The specific task instances we utilised were:
The types of arenas that these instances correspond to can be found at (http://animalai.org/AAI/testbed).
Appendix C AAIO Agent Results
This section gives more details for the first domain we examined in Section 4, the Animal-AI Olympics competition dataset. We also give more detail about the specifics of the measurement layouts.
C.1 Patterns of Performance
To show the patterns of performance of several agents, we use a simple visualisation known as agent characteristic grids, as described in (Burnell et al. 2022). The plots place each feature as a dimension and the success rate represented in colours from green (success) to red (failure). Values are grouped into bins of appropriate width in the desired dimensions for analysis. Each cell also shows the number of observed instances in that bin. Fig. 8 shows some of the agents discussed in the paper, as well as some other agents that will serve us to illustrate a few points.
For instance, ‘ironbar’ has 99% accuracy (the grid is mostly green everywhere) for this set of tasks (like ‘Trrrr’, the winner of the competition) and is identified as saturating both capabilities in Fig. 4. The next one, ‘sparklemotion’, with 36% accuracy, is very non-conformant, with navigationAbility $0.40\pm 0.38$ and visualAbility $0.59\pm 0.45$, and high noise $0.31\pm 0.13$. Proportionally, navigationAbility is worse as the grid looks even less conformant on the distance of the reward than the reward size. The rightmost on the top row is ‘Juohmaru’, with 54% accuracy, which is a very conformant agent, with red concentrating on the top and right of the grid (the objectively most difficult instances). Its navigation ability is $2.50\pm 1.08$ and visual ability is $0.92\pm 0.44$, and medium noise: $0.26\pm 0.17$.
C.2 Animal AI Olympics Measurement Layout
Table 2 provides more information about the task meta-features that we utilised to construct the measurement layouts. Note that these ranges determine the ranges of the capabilities, so that we have units. For instance, if distance to the reward were measured in metres then the navigation capability would have metres as a unit. Let the navigation capability of an agent be, say, 20 metres. Since it is defined with a logistic function of distance, that means that the agent can achieve excellent performance when well below 20, performing at chance around the 20 metre mark, and failing every instance when the reward is well beyond 20 metres away.
Dimension
Description
Range/Values
Reward Size
The size of the reward (the sign is modified so lower is larger)
$[0,1.9]$
Reward Distance
Euclidean distance from the agent’s start position to the reward
$[0,5.3]$
Reward Behind
Whether the reward originates behind (1), left or right (0.5) or in front (0) of the agent
$\{0,0.5,1\}$
XPos
Whether the reward is to the left ($-1$), the right (1) or is centred (0) relative to the agent’s start position
$\{-1,0,1\}$
Name Type Range/Values Prior / Linking Function
rewardSize Meta-Feature $[0,1.9]$
rewardDistance Meta-Feature $[0,5.3]$
rewardBehind Meta-Feature $\{0,0.5,1\}$
XPos Meta-Feature $\{-1,0,1\}$
navigationAbility Capability $[0,5.3]$ $ScaledBeta(1,1,0,5.3)$
visualAbility Capability $[0,1.9]$ $ScaledBeta(1,1,0,1.9)$
rightleftBias Bias [,][$\infty$,$\infty$][ - ∞ , ∞ ] $\mathcal{N}(0,1)$
noiseLevel Robustness $[0,1]$ $Uniform(0,1)$
rightleftEffect INON $[-1,1]$ ${\sf rightleftBias}\times{\sf XPos}$
Table 3: Nodes in the Animal AI Measurement Layout, including range/values, priors and linking function.
The measurement layout for this suite of tasks consists of two capability nodes (navigation and visual acuity), one bias node (whether the goal is initially on the left or the right of the agent), and one robustness node (stochastic noise on success/failure). Each of the capability and bias nodes has an associated instance-performance node, its intermediate non-observable node (INON). The noise is applied at the end.
In Table 3 we have complete details of all the nodes in the measurement layout. The function $\sigma$ represents the standard logistic function:
$\sigma(x)=\frac{1}{1+e^{-x}}$
while $\omega$ represents a weighting:
$\omega[\alpha](b,c)=(1-\alpha)\cdot b+\alpha\cdot c$
The value $\nu$ represents a noise prior that represents a constant model set to 1 -- average performance. Finally, $ScaledBeta$ represents a scaled beta distribution to an interval different from $[0,1]$. Note that a beta distribution has two parameters that need to be estimated.
Figure 9: Inferred cognitive profile elements, along with the success rate, for the 68 agents in the AAIO. The gradient colour red-yellow-green denotes how high the means are (green is highest) while the gradient circles white-grey-black denote how high the standard deviations are (black is highest).
Finally, Fig. 9 shows the complete inference results for the 68 agents in the AAIO. The agents are ordered by the aggregate success rate at the collection of task instances. We see that, for the agents in the middle of the table in particular, our approach is able to decompose agent capabilities at this task into the constituent navigation and visual abilities. Bias has a more diverse behaviour throughout the table, with strong biases for extremely low scoring agents too, such as with agents ‘ahorizon’ and ‘Octopus’.
Appendix D O-PIAAGETS Battery
Now we look at the object permanence test battery in more detail. We look at the overall design criteria for the test-bed, the types of tasks involved, and how the tasks are characterised. We also give full details for the measurement layout devised for this task and an extended view of the results inferred by our approach.
D.1 Design Criteria
Imagine a simple task where a reward is observed to roll leftwards behind an occluder, and go fully out of sight. An agent with OP ought to pass this task, since it: (a) wants to obtain the reward, (b) understands that the reward continues to exist despite not being able to see it, and (c) is able to infer the location of the occluded reward based on its previously observed trajectory. However, an agent with OP might fail this task, if for instance it is bad at navigating, and takes a very circuitous route to its goal and runs out of time before it can reach the reward. Conversely, an agent lacking OP might pass this task, simply because it has been rewarded previously for going forwards some number of steps and then turning left. This is the problem of the underdetermination of theory by behaviour (Stanford 2006), in which an observed behaviour can be explained in multiple ways. The use of well-informed measurement layouts based on internally valid test batteries allows us to more robustly eliminate incorrect alternative explanations and triangulate on the underlying causal explanations of AI system behaviour.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
These considerations were used as design criteria for the test-bed presented in the next subsection.
D.2 Object Permanence Tasks
The tasks that we use in our OP test-bed are based on a subset of the O-PIAAGETS object permanence test battery introduced in (Voudouris et al. 2022). The battery in its entirety is vast (including over 13,000 task instances requiring object permanence and over 250,000 basic task instances that don’t require OP, but help control for other potential explanations. Here, we adapt the following three separate task types from O-PIAAGETS.
Primate Cognition Test Battery Cup Tasks
Herrmann et al. developed the Primate Cognition Test Battery (PCTB) for testing aspects of physical and psychological common-sense in primates (Herrmann et al. 2007). One task in this battery was designed to assess OP. The participant is presented with three identical inverted cups, under which the experimenter hides a reward. Once the reward has been occluded, the participant is invited to lift a cup. If they fail to locate the reward, the trial ends. The reasoning here is that if the participant correctly locates the reward without any evidence of trial-and-error, there is evidence that they have OP, understanding that the reward continues to exist even when it is out of view.
Figure 10: Primate Cognition Test Battery 3-Cup (’Cup’) Task. Top: Top-down view; bottom: Agent-view. Red arrow indicate reward location, grey arrows indicate agent location.
O-PIAAGETS has a battery of 598 instances based on this experimental set-up, including 359 that require object permanence to be solved and 239 that do not. In the OP tests, the agent observes a green reward drop from a height behind an occluder. They must traverse a ramp, sometimes avoiding lava, to obtain it. If they traverse the wrong ramp, they cannot return, and they will fail the instance, making this a forced choice test. See Fig. 10. In the instances that do not require OP, the reward is always visible, either through transparent objects or due to the lack of occluders. These instances serve as controls, so that we can evaluate whether poor performance on tasks requiring OP can be explained not by a lack of object permanence per se, but rather by poor performance traversing ramps, avoiding lava, or navigating towards rewards.
Primate Cognition Test Battery Grid Tasks
Extending the general structure for testing OP, Crosby et al. developed a task where rewards can be occluded in one of up to 12 different positions, through holes in the floor (Crosby et al. 2020). O-PIAAGETS contains 430 instances based on this design, with 4, 8, or 12 holes (see Fig. 11). 191 instances require OP to be solved, with the remaining 239 not requiring OP, where the holes are not deep enough to occlude the goal. These latter instances serve as controls.
Figure 11: Primate Cognition Test Battery Grid (’Grid’) tasks. Top: Top-down view; bottom: Agent-view. Red arrow indicate reward location, grey arrows indicate agent location. Reproduced from Voudouris et al., 2022.
Chiandetti & Vallortigara Chick Tasks
Chiandetti and Vallortigara developed a series of tests for investigating physical common-sense in day-old chicks (Chiandetti and Vallortigara 2011), which Crosby et al. developed for use in the Animal-AI Environment (Crosby et al. 2020). In the example in Fig. 12, the agent observes a reward rolling away (panel A). Then the lights go out, blocking all visual input for several time steps (panel B). During this lights out period, the reward is deflected either left and rolls behind the occluder. When the lights turn on, the reward is fully occluded (panel C). If the agent understands that the reward continues to exist even when it is out of view, they ought to reason that it is behind the occluder on the left, as there is nowhere it could be hiding on the right.
Figure 12: Chiandetti and Vallortigara (’CV’) Task. Top: Reward Rolling away; middle: ”Lights out” period; bottom: Reward occluded. Reproduced from Voudouris et al., 2022
We used 36 instances from O-PIAAGETS that tested OP in this way, and included 1195 further tasks that did not require OP to be solved, to serve as controls.
Basic Tasks
All of these tasks require an agent to be able to navigate towards green goals, around obstacles, while traversing platforms and ramps, and avoiding lava. When testing for OP in biological agents such as primates or chicks, it is often safe to assume that they understand the value of rewards such as food or that they are able to navigate or search for rewarding objects. This is not a safe assumption when evaluating artificial agents, particularly those trained purely with techniques from machine learning. Therefore, the testbed must include instances that test for these basic skills. We included a further 521 instances testing basic abilities to navigate towards reward, avoid lava, and use ramps. These are extensions and variations of the tests used in section  4.
In total, we included 2188 unique instances from O-PIAAGETS, varying along the 12 dimensions outlined in Table  4. Note the asterisks in the table, denoting that a particular feature was not used in the measurement layouts used for evaluation in this paper. This occurred because Team B either didn’t realise the feature was present in the environment or because it was missed as a possible source of bias or capability. Nevertheless, this actually highlights a strength of our evaluation framework; imperfect measurement layouts can still be useful for both inferring agent capabilities and predicting future agent performance.
Dimension
Description
Range/Values
Reward Visibility*
Is the reward visible at the choice point?
${0,1}$
Time Under Occlusion
Time reward is occluded
$[0,4]$
Positions
Number of alternative reward places
$[1,11]$
Reward Distance
Manhattan distance to the goal
$[0,56]$
Right/left Position
Is reward left, right or straight from the agent?
$\{-1,0,1\}$
Reward Size
How large is the reward?
$[0.5,4]$
Occluder Presence
Are there any occluders hiding the rewards?
$\{0,1\}$
Lava Presence
Is there any lava in the arena?
${0,1}$
Platform Presence
Are there any platforms in the arena?
${0,1}$
Ramp Presence
Are there any ramps in the arena?
${0,1}$
Transparent Walls
Are there any transparent walls in the arena?
${0,1}$
Occluder Redness*
What is the R value of occluders in the arena?
$[0,255]$
Occluder Greenness*
What is the G value of occluders in the arena?
$[0,255]$
Occluder Blueness*
What is the B value of occluders in the arena?
$[0,255]$
Table 4: Object Permanence Task Characterisation details. The dimensions that instances vary on. The dimensions with an asterisk are not included in the measurement layout presented in this paper.
This test-bed satisfies our criteria for capturing and evaluating OP that we outlined above:
  1. 1.
    Object permanence is difficult to place on an ordinal scale. In cognitive science, it is usually discussed in terms of its presence or absence.
  2. 2.
    Object permanence relies on working memory capacity and spatial cognition, and successful performance on tests of OP often require good navigational skills. The dimensions defining each task vary on their relevance to OP and these precursor abilities (working memory, spatial cognition, navigational ability). There is therefore a complex interaction between latent capabilities and task dimensions.
  3. 3.
    Instances were not defined on all dimensions, introducing missingness. For example, in many instances, there were no occluders, so Occluder Colour is undefined.
  4. 4.
    Due to the structure of Animal-AI and O-PIAAGETS, some instance-types are represented more than others. For example, we have 36 Chiandetti & Vallortigara Chick Tasks testing OP, compared to 359 Primate Cognition Test Battery Cup Tasks, almost 10 times more.
D.3 Object Permanence Measurement Layout
This section extends the information about the measurement layout in section 5.2 and Fig. 5. The measurement layout for OP consists of seven capability nodes, one bias node, and one robustness node (stochastic noise on success/failure). As in the AAIO case, each of the capability and bias nodes has an associated instance-performance node, an intermediate non-observable node (INON). The noise is applied at the end to the final performance probability.

### v2

@@ C.2 Experiment 2: An Object Permanence Task
Team A generated 30 synthetic agents of four different types to ensure a diversity of behaviour. Reference Agents act as reference points, e.g., a perfect agent, an agent with xx% success for various values of $x$.
@@
Fraudster Agents behave in a manner that might look like object permanence, but are actually using rigid heuristics such as “go to last seen location of reward, then search” (Užgiris and Hunt, 1975). Tables  5,  6, and 7 in Appendix E.1 show the specifics of each agent. This information was not disclosed to Team B.
C.3 Experiment 3: Extending Measurement Layouts To Real Data
We use the dataset from Voudouris et al. (2024) to fit measurement layouts for realistic data collected from deep reinforcement learning agents and human children, as well as a heuristic agent and a random action agent.
We use the random agent to define a baseline for chance performance. This agent takes one of the nine available actions available in the Animal-AI Environment and repeats it for a number of steps, $S$ sampled from a uniform distribution: $S\sim U(0,20)$, after which it randomly selects a new action. It executes an action for that many steps before selecting a new one. There are no biases to select certain actions.
We use the heuristic agent as a more meaningful baseline, as an agent that lacks object permanence, but possesses many other capabilities that would lead to success in the Animal-AI Environment. We do not expect random walkers to fail tests of OP because they lack only the capability of OP. Rather, we expect them to fail because they lack many other capabilities, such as the capability to move towards an appetitive reward or away from an aversive stimulus (goal-directedness). The heuristic agent follows a small set of rules and heuristics. This agent can view objects in the $x$-$z$ plane that are in the forward-facing 60° of its view using raycasts. When it detects a green or a yellow goal in this viewing angle, it orients itself towards it and moves forward to obtain it. When it detects a red goal or a death zone, it orients itself away from it and moves backwards. If the agent is stationary or if there is a wall in front of it, the agent moves either forwards and left or forwards and right with a probability pp=0.5, perseverating that movement subsequently to navigate around the object until the agent is no longer stationary or there is no longer a wall in front of it. If the agent does not detect any goals, death zones, or walls, and if it is not stationary, it perseverates its previous action.
Data from children aged 4–7 years old performing on a subset of the O-PIAAGETS tasks was also used as a meaningful baseline for an agent that is error-prone but possesses object permanence and other ancillary capabilities necessary to complete tasks in Animal-AI Baillargeon et al. (2010b); Užgiris and Hunt (1975); Piaget (1923). Data from children was collected between January and March (inclusive) 2023 by the authors of Voudouris et al. (2024), with a minimum target sample size of 30. The study was conducted at a university, in the presence of researchers and the participants’ guardians. Guardians were provided with an information sheet and the opportunity to consent to their child’s participation in the study. If they consented, the study was initiated. Participants watched a three-minute video introducing the study as a game called ‘Get the Fruit!’. Participants were invited to assist ‘Farmer John’ in finding all the green and yellow ‘apples’ in the farmyard, while avoiding any poisonous red fruit and ‘large puddles of lava rain.’ They were informed that they would receive stickers in return for fruit collection, although they actually received one sticker every ten trials, regardless of whether they successfully found the goal, and an extra sticker upon completion of the study. After the video, participants commenced play with the Animal-AI Environment, interacting with it through a small hand-held controller connected to a computer. First, they played a ‘tutorial’ round, consisting of 11 tasks. They could play these tasks as many times as they wished, and ask for help throughout. Then, they played the ‘test’ round, consisting of 38 Test tasks. These tasks were ordered randomly, with half of participants playing the fixed random order and the other half playing the reverse of that order. Every 10 tasks, the participant could take a break from the game and receive their stickers. The break consisted of a simple task with a single large green goal and no time limit. When the participant was ready to continue playing, they simply had to navigate to the goal to initiate the next batch of tasks. During the test round, guardians filled out a short survey about the participant, with questions on age, gender, and videogame playing habits. Upon completion of the test round, participants and their guardians were debriefed and offered a book voucher for their participation.
The deep reinforcement learning agents, Dreamer-v3 and PPO, were each trained on 5 different curricula, resulting in 5 different trained agents (see Table 2).
Table 2: The 5 different trained agents for each architecture (Dreamer-v3 & PPO; 10 total), their training curricula, and the number of training steps. Stratification was done by experimental paradigm.
Agent Name Training Curriculum Training Steps (Millions)
PPO/Dreamer 1 All Basic tasks 2M
PPO/Dreamer 2 All Basic tasks ++ a stratified random sample of 300 Control tasks 4M
PPO/Dreamer 3 All Basic tasks ++ all Control tasks (in 3 randomly sampled batches) 8M
PPO/Dreamer 4 All Basic tasks ++ a stratified random sample of 300 Control tasks ++ a stratified random sample of 300 Test tasks 6M
PPO/Dreamer 5 All Basic tasks ++ all Control tasks (in 3 randomly sampled batches) ++ a stratified random sample of 300 Test tasks 10M
Appendix D Measurement Layout Specifications
D.1 Experiment 1: A Simple Navigation Task
The measurement layout for the simple navigation task consists of two capability nodes (navigation, visual acuity) and two meta-feature nodes (goal distance, goal size), as depicted in Figure 2B. Each of the capability nodes has an associated instance-performance node, an intermediate non-observable node (INON). To compute probabilities at these nodes, we use the sigmoid function over the difference between estimated capability and the current meta-feature value. For visual acuity, this is the goal distance divided by the goal size. We apply the natural logarithm to this difference to improve numerical stability. For navigation, the meta-feature is simply the goal distance. We assume a non-compensatory relationship between navigation and visual acuity, and therefore multiply these marginal INON probabilities to condition the Bernoulli distribution over binary (success/failure) outputs. We use the half-normal prior over capabilities with a $\sigma$ equal to half of the maximum relevant meta-feature or meta-feature combination.
D.2 Experiment 2: An Object Permanence Task
The measurement layout for OP consists of seven capability nodes, one bias node, and one robustness node (stochastic noise on success/failure). Again, each of the capability and bias nodes has an associated instance-performance node, an intermediate non-observable node (INON). The noise is applied at the end to the final performance probability. A full description is given in Table 3 including all the value ranges, distributional priors and linking functions.